# Critical modular lattices in the Gaussian core model

### v1

@@ 5.6. A brief remark on simultaneous diagonalizability
[…]
@@
hence, $T$ is an $\operatorname{O}(L)$-equivariant endomorphism of $\mathcal{S}^{n}$. In other words, $T$ is intertwining the representation $(\mathcal{S}^{n},\langle\cdot,\cdot\rangle)$ of the group $\operatorname{O}(L)$ with itself. In this way we have a $1$-$1$-correspondence between $\operatorname{O}(L)$-invariant bilinear forms $Q$ on $\mathcal{S}^{n}$ and $\operatorname{O}(L)$-equivariant endomorphisms of $\mathcal{S}^{n}$.
@@ 6. Outlook
The strategy described in Section 3 and carried out for some examples in Section 5 can likely be applied to other extremal (strongly) $\ell$-modular lattices as discussed in [23], [24], and [27]. However, several adjustments have to be made: Firstly, one needs explicit information about the spaces $\mathcal{S}_{n/2+4}(\Gamma_{0}(\ell,\chi))$, in particular if $\chi$ is non-trivial, to write down the data needed for an application of Theorem 3.3 (or a variant of it). This includes the dimension of this space and a basis as in (19), or it would involve more complicated computations following (18). Next, one needs explicit bounds on the coefficients of the cusp forms in $\mathcal{S}_{k}(\Gamma_{0}(\ell),\chi)$, while we are aware of such bounds only in the cases $\mathcal{S}_{k}(\Gamma_{0}(2))$ and $\mathcal{S}_{k}(\Gamma_{0}(3))$. Analogous bounds for larger $\ell$ and characters $\chi$ would allow for rigorous computations as presented throughout Section 5.

### v2

@@ 5.8. A brief remark on simultaneous diagonalizability
This turns $(\mathcal{S}^{n},\langle\cdot,\cdot\rangle)$ into a unitary representation of $\operatorname{O}(L)$, meaning that the action of $\operatorname{O}(L)$ preserves the inner product $\langle\cdot,\cdot\rangle$. By virtue of the Riesz representation theorem, there is a linear map $T:\mathcal{S}^{n}\to\mathcal{S}^{n}$ such that
@@
hence, $T$ is an $\operatorname{O}(L)$-equivariant endomorphism of $\mathcal{S}^{n}$. In other words, $T$ is intertwining the representation $(\mathcal{S}^{n},\langle\cdot,\cdot\rangle)$ of the group $\operatorname{O}(L)$ with itself. In this way we have a 11-11-correspondence between $\operatorname{O}(L)$-invariant bilinear forms $Q$ on $\mathcal{S}^{n}$ and $\operatorname{O}(L)$-equivariant endomorphisms of $\mathcal{S}^{n}$.
@@ 6. Outlook
The strategy described in Section 3 and carried out for some examples in Section 5 can likely be applied to other extremal (strongly) $\ell$-modular lattices as discussed in [27], [28], and [31]. However, some adjustments have to be made: Firstly, one needs explicit information about the spaces $\mathcal{S}_{n/2+4}(\Gamma_{0}(\ell,\chi))$, in particular if $\chi$ is non-trivial, to write down the data needed for an application of Theorem 3.3 (or a variant of it). This includes the dimension of this space and a basis as in (19), or it would involve more complicated computations following (18). Next, one needs to carry out computations to bound the coefficients of the cusp forms in $\mathcal{S}_{k}(\Gamma_{0}(\ell),\chi)$, for which one can use the method described in Section 4.2, if one can find a decomposition into Hecke eigenforms. Analogous bounds for larger $\ell$ and characters $\chi$ would allow for rigorous computations as presented throughout Section 5.
Acknowledgements
We thank Matthew de Courcy-Ireland and the anonymous referee for pointing out how to bound the coefficients of cusp forms using the bound on Hecke eigenforms coming from Deligne’s theorem, improving the resulting bounds by certain orders of magnitude compared to the previously used global bounds.
We further thank the anonymous referee for the helpful comments, suggestions, and corrections on the manuscript.
We thank Henry Cohn for pointing out an error in an earlier version.